# Evidence of equilibrium dynamics in human social networks evolving in time

### v1

@@ Abstract
The dynamics of personal relationships remain largely unexplored due to the inherent difficulties of the longitudinal data collection process. In this paper, we analyse a dataset tracking the temporal evolution of a network of personal relationships among 900 people over the course of four years. We search for evidence that the network is in equilibrium, meaning that all macroscopic properties remain constant, fluctuating around stable values, while the internal microscopic dynamics are active. We find that the probabilities governing the network dynamics are stationary over time and that the degree distributions, as well as edge and triangle abundances match the theoretical equilibrium distributions expected under these dynamics. Furthermore, we verify that the system satisfies the detailed balance condition, with only minor point deviations, confirming that it is indeed in equilibrium. Remarkably, this equilibrium persists despite a high turnover in network composition, suggesting that it is an inherent characteristic of human social interactions rather than a trait of the individuals themselves. We argue that this equilibrium may be a general feature of human social networks arising from the competition between different dynamical mechanisms and also from the cognitive and material resources management of individuals. From a practical perspective, the fact that networks are in equilibrium could simplify data collection processes, validate the use of cross-sectional data-based methods like Exponential Random Graph Models, and inform the design of interventions. Our findings advance the understanding of collective human behaviour predictability and our ability to describe it using simple mathematical models.
@@ 1 Introduction
Human social behaviour is driven by a complex interplay of cognitive, emotional, and social factors, which together shape the structure and dynamics of social networks. Focusing on personal relationships, prior research has shown that individuals tend to maintain a finite number of social bonds constrained by their cognitive capacity [1], and how these social bonds are structured is explained by the way humans allocate their cognitive resources to create functional relationships able to cover their social necessities [2, 3]. From a dynamical point of view, individuals continually adjust their social ties in response to emotional, environmental, and cognitive changes. Still, this dynamic behaviour of social bonds is less understood, mainly due to a lack of rich longitudinal data: longitudinal studies on social networks evolution are rare and often face issues like small sample sizes, noisy data, and limited conclusions. Our work focuses on the temporal evolution of social relationships and addresses the question of whether these complex tendencies give rise to stable trends.
Due to the interconnected nature of social ties, network theory constitutes an appropriate framework to investigate such phenomena. In network science [4], and in particular in the context of social networks [5], personal relationship networks are defined as networks where nodes represent people and links represent personal relationships, such as friendships or enmities. In general, data on these networks is gathered by selecting a sample population of unrelated individuals and asking each person to name the people with whom they have a relationship [6] and how they are related, or by surveying individuals within a closed environment, such as a school, university, or company [7]. Examples of datasets collected using the first strategy include the Caen Panel Survey [8], analysed extensively in [9, 10, 11, 12], and other similar datasets replicated on a smaller scale, like [13]. Regarding the second strategy, important datasets include Newcomb’s Fraternity data [14, 15, 16, 17, 18], and Sampson’s data [19]. Several studies have also collected data in school environments [20, 21, 22, 23, 24, 25], being specially relevant the National Longitudinal Study of Adolescent to Adult Health (Add Health) [26, 27]. These datasets represent an outstanding effort to overcome the challenges of collecting longitudinal data. However, they have limitations that constrain the conclusions that can be drawn from their analysis, aside from small sample sizes (with the exceptions of [24, 27]). The first method’s data are usually undirectional and binary, as the respondent reports the mere existence of third-party relationships according to their perception. The second method provides directed but still binary and partial data, often limited to within-class relationships with an artificial cut-off (of typically 5) to the number of relationships reported by the respondent imposed by the surveyor. Furthermore, many studies collected only two to three snapshots, limiting long-term dynamic analysis.
Given these limitations, some alternative ways to investigate the temporal evolution of personal relationships have been attempted. A possible strategy involves inferring the network of personal relationships from indirect data, such as phone calls [28], email exchanges [29] or face-to-face interaction data [30, 31], instead of using surveys. The only drawback is that conclusions about personal networks depend on the assumption that this indirect data is indeed a proxy of the underlying personal relationships. Other researchers have chosen to study network dynamics indirectly using cross-sectional data, which is easier to obtain. The basic idea behind this approach is that the dynamical mechanisms driving the temporal evolution of a network leave a fingerprint in the structure observed at single points in time. For instance, if nodes tend to reciprocate links, a static network snapshot will show a higher number of reciprocal links than expected by chance. This is the key idea behind Exponential Random Graph Models [32], which infer network formation mechanisms based on the observed structure, assuming that the observed network is a random sample from a family of networks produced by the model, and behind other similar methods [33]. Finally, a third strand of research has been more focused on developing models with specific hypothetical mechanisms and simulating the network’s evolution according to these models [34, 35, 36]. Closely related, some works make use of traditional statistical mechanics of the equilibrium techniques to explore structural patterns in networks [37, 38].
@@
Using these data and methods, several studies have focused on specific aspects of the evolution of social ties. For instance, Rivera et al. explore the temporal dynamics of dyadic relationships [39], while Yap and Harrigan assess how various mechanisms can explain the observed behaviour across different datasets [40]. In contrast, our work takes a broader approach, focusing on the collective behaviour of social relationships. Specifically, we analyse whether personal relationship networks can exhibit stable trends over time, a phenomenon we refer to as equilibrium dynamics. In the following sections, we will provide a rigourous definition of this concept inherited from statistical physics. In short, we consider a social network to be in equilibrium when its macroscopic, average properties – such as the distribution of personal contacts; the prevalence of patterns like edges, triangles, or larger structures; and aggregate metrics like network density – remain constant, fluctuating around stable values, even as individual links continuously evolve at a microscopic level. Our interest is driven by the lack of a rigourous exploration within the framework of equilibrium dynamics of some of the mentioned studies [20, 23, 17, 25, 29, 28, 13, 41]. In particular, and interestingly, there are apparent trends of stability in other aspects of human social behaviour, which suggest a natural tendency toward the formation of stable patterns and self-reinforcing dynamics. For example, Hobbs and Burke show how social connections recover after the death of a friend [42], while Alessandreti et al. found that certain mobility patterns are conserved over time [43], linking these patterns to individuals’ social ties. Other studies, such as [44, 45, 46], stress the stability and robustness of online social behaviours. We will circle back to this research in the discussion section.
The idea that human social networks may be in equilibrium is relevant from the perspective of social behaviour because it suggests that, despite the complexity of individual interactions and the interplay between cognitive, emotional, and environmental factors, the overall structure of these networks remains stable over time. This stability points to a fundamental predictability in social systems, implying that, like physical systems, human networks follow certain rules, and their evolution is constrained. We will discuss how these constraints may arise from cognitive and material resource management, with individuals regulating their social ties to avoid overwhelming their social capacities. Besides, if a social network is in equilibrium, its properties can be predicted more accurately over time, facilitating intervention studies. Interventions can be introduced and compared against the equilibrium network, which serves as a baseline for analysis, helping to strengthen social cohesion or address social isolation. The possibility that this equilibrium behaviour may be general suggests that certain social behaviours and organization principles transcend cultural, situational or individual differences, connecting individual interactions and broader social structures. From a modelling perspective, equilibrium assumptions enable the use of simpler, more tractable mathematical models, such as statistical mechanics tools, developed for physical systems.
From a practical point of view, if social networks are in equilibrium, the potential for generalizability and robustness of findings from different studies significantly increases. Conclusions about the mechanisms driving network dynamics, structural trends, and other insights apply not only to a single observed point in time but to the entire unobserved evolution of the network after the transient formation period. This is specially relevant for methods that employ cross-sectional data to investigate network dynamics, like Exponential Random Graph Models. These approaches have an implicit and often overlooked assumption: the network must be in or near equilibrium, as discussed in [47]. In simple terms, when using cross-sectional data, the researcher assumes that every observation is statistically equivalent to any other, a condition that is fulfilled if the network is in equilibrium. This reduces the need for continuous data collection, lowering resource requirements and the burden on respondents.
Despite its importance, to the best of our knowledge, no prior research has explicitly focused on determining whether social networks are indeed in equilibrium. In this context, we present an example of an empirical social network that is in rigourous equilibrium, given by a rich dataset collected over four years. We also explore the implications of this finding, and, we discuss whether this should be a general phenomenon in any social network.
@@ 2 Results
In this work, we analyse a dataset we have collected between 2020 and 2024 that contains the temporal evolution of the network of personal relationships among 888 people belonging to the Blas de Otero High School in Madrid (see the Materials and Methods section for details on the data collection, data composition and data curation). The reported relationships were coded as +2 (very good), +1 (good), -1 (bad), and -2 (very bad). Almost every student provided this list, resulting in the extraction of a weighted directed network for the entire high school, which we will refer to as snapshot or wave. We repeated this process every 16-20 weeks, and with this data we reconstructed the network at different points in time, collecting 10 snapshots of the network over 4 years. With respect to previous literature, we have introduced some improvements in the data collection process. Respondents are allowed to report an unlimited number of relationships, resulting in richer and more heterogeneous data. We also increased the sample size by one order of magnitude, with a larger number of snapshots, allowing for the detailed study of the dynamics in the long term while maintaining a certain level of granularity. Relationships are self-reported, and they are no longer binary. Instead, we include directions and weights to the relationships, between -2 and 2, allowing the consideration of negative relationships. Preliminary analyses of earlier, shorter versions of this dataset have been reported elsewhere [48, 41]. It is worth noting that by coding relationships into these four categories (-2, -1, +1 and +2), we inevitably lose many of the details that characterize social interactions. Thus, the resulting network we analyse in this study is not a direct reflection of the complexity of real personal relationships, but rather a simplified representation. Real-world relationships exist on a spectrum, with layers of emotional, social, and contextual subtleties that cannot be completely captured in these discrete categories. However, even with this simplification, we argue that identifying an equilibrium in this network serves as a meaningful proxy for the underlying dynamics of the real network. The equilibrium we detect in this reduced representation suggests that the macroscopic patterns of stability we observe likely reflect deeper robust social structures. If the simplified relationships follow this equilibrium dynamics, it implies that the more complex relationships they represent also tend toward stability over time.
@@
Drawing an analogy to a social system represented as a network, it is essential to define both the macroscopic properties and the microscopic dynamics. In a social network evolving over time, links appear, disappear, or change in nature. From the perspective of a node, an outgoing or incoming +1 link can become a +2, a -1, a -2 or disappear. We refer to an absent link as a 0 link. This constitutes the microscopic dynamics of the system: the evolution of individual ties/links. From these microscopic components, we can build macroscopic properties of the network. In the network, we define an edge as the connection between two nodes that contains two links, one from the first node to the second and one from the second node to the first. Since we have 5 types of links, -2, -1, 0, +1, +2, there are 25 possible edge types. Notice the distinction between link/tie and edge we introduce in our notation. Although it is usual to treat both as interchangeable, we keep this distinction throughout the paper. Our first macroscopic property is the distribution of different edge types: how many +2+2, +1+2, +1+1, etc., edges exist in the network. Notice that a +2+1 edge is not equivalent to a +1+2, specially from a dynamical perspective. Although both edges can evolve towards a +1-1 edge, in the +2+1 case the +2 needs to become a +1 and the +1 a -1, and in the +1+2 case, the +1 remains unchanged and the +2 becomes a -1. Therefore, we keep this distinction in all computations. A second macroscopic property is the in-degree and out-degree distributions of the nodes. For each type of link (+2, +1, -1, and -2), we count how many nodes have 0 incoming +2 links, 1 incoming +2 links, etc. This process is repeated for outgoing links and every type of link, generating eight degree distributions. A third macroscopic property can be the distribution of different triangle types that can form among three nodes. Other macroscopic properties include structural characteristics of the network, such as clustering, average shortest path length, assortativity, etc.
@@
Let us introduce the concept of a transition matrix using the edges as our macroscopic property. In the case of edges, there are 25 possible edge states. If we take two consecutive snapshots of the network, for each individual edge, we can record what kind of transition it went through. A +2+2 edge can stay a +2+2, or become a +2+1, or even a +2-1, etc. If we repeat this process for all the links, we can construct a 25x25 matrix in which each row is the edge state before the transition, in the first snapshot, and each column is the edge state after the transition, in the second snapshot. The $ij$ element of this matrix would be the number of edges starting in state $i$ and ending in state $j$. In this matrix, we can divide each element by the total count in each row. By doing this, the $ij$ element in this final matrix represents $P(j|i)$, i. e., the conditional probability of an edge of going to the $j$ state provided it started in the $i$ state. We call this matrix the transition matrix $P_{\beta\alpha}$ (also known as the Markov Matrix), from snapshot $\alpha$ to snapshot $\beta$. Additionally, we define the edge state distribution $\pi_{\alpha}$ as a column vector in which each element $i$ is the density of edges of type $i$ in the snapshot $\alpha$. With these two objects, it is direct to see that:
@@
In other words, if we take the distribution of edges in one snapshot and multiply this distribution by the transition matrix, we obtain the distribution of edges in the next snapshot.
Since we have 10 snapshots, our data allows us to construct nine transition matrices between consecutive snapshots. The first relevant question is whether these nine transition matrices are statistically equivalent. If they were, it would indicate that the dynamics are stationary, meaning the transition probabilities between edge states remain constant over time. However, notice this does not necessarily imply that the system is in equilibrium. We could have stationary dynamics where macroscopic properties change over time, but these changes would be uniform because the transition probabilities do not vary.
To simplify the process, we check whether the nine transition matrices are statistically equivalent to the average transition matrix computed by averaging the nine individual matrices. To compare each transition matrix with the average, we resampled the individual matrix randomly from the data 1000 times. We calculated the distance between each element of the randomly sampled matrix and the corresponding element in the average transition matrix. To do such resampling, we created a pool with all the observed transitions for each individual edge in each pair of waves, removing the information about the actual waves in which that individual transition occurred. We then created each randomly resampled matrix by drawing samples at random from this pool with no information about the specific waves in which transitions occurred. We counted how often each element of the resampled matrix was farther from the corresponding element in the average matrix than the actual value for the matrix we wanted to compare. In summary, we calculated, element by element, the probability that a randomly sampled matrix element would be farther from the corresponding element of the average matrix than the same element in the actual transition matrix we were analysing.
@@
It is worth noting that we perform an element-wise comparison. Although we could compare the matrices themselves using metrics like the Frobenius distance, we chose a granular approach to be able to identify individual differences. However, it is important to take into account that this method involves multiple comparisons of elements in a sampled matrix; specifically, we are comparing 625 elements against another 625 elements. Consequently, we expect some elements to deviate more than expected by chance, even if the matrices are equivalent. For such cases, we compute a z-score to determine how significant these deviations are from randomness.
@@
As mentioned, the stationarity of the transition matrices does not necessarily imply that the network is in equilibrium; it may not even have an equilibrium state under these dynamics. As we will see, this is not our case.

### v2

@@ Abstract
How do networks of relationships evolve over time? We analyse a dataset tracking the social interactions of 900 individuals over four years. Despite continuous shifts in individual relationships, the macroscopic structural properties of the network remain stable, fluctuating within predictable bounds. We connect this stability to the concept of equilibrium in statistical physics. Specifically, we demonstrate that the probabilities governing network dynamics are stationary over time, and key features like degree, edge, and triangle abundances align with theoretical predictions from equilibrium dynamics. Moreover, the dynamics satisfies the detailed balance condition. Remarkably, equilibrium persists despite constant turnover as people join, leave, and change connections. This suggests that equilibrium arises not from specific individuals but from the balancing act of human needs, cognitive limits, and social pressures. Practically, this equilibrium simplifies data collection, supports methods relying on single network snapshots (like Exponential Random Graph Models), and aids in designing interventions for social challenges. Theoretically, it offers new insights into collective human behaviour, revealing how emergent properties of complex social systems can be captured by simple mathematical models.
@@ 1 Introduction
Human social behaviour is driven by a complex interplay of cognitive, emotional, and social factors, which together shape the structure and dynamics of social networks. Prior research has shown that individuals tend to maintain a finite number of social bonds constrained by their cognitive capacity [1], and how these social bonds are structured is explained by the way humans allocate their cognitive resources to create functional relationships able to cover their social necessities [2, 3]. From a dynamical point of view, individuals continually adjust their social ties in response to emotional, environmental, and cognitive changes. Our work goes beyond the temporal evolution of personal social relationships, focusing on whether these complex tendencies give rise to stable trends in the emergent social structure.
In the context of social networks [4], personal relationship networks are defined by nodes representing people and links representing personal relationships, such as friendships or enmities. Previous research has studied the dynamics of these personal networks, specially using data collected in closed environment through surveys. Examples include the Newcomb’s Fraternity data [5, 6, 7, 8, 9], and Sampson’s data [10]. Other studies have collected data in school environments [11, 12, 13, 14, 15, 16], the National Longitudinal Study of Adolescent to Adult Health (Add Health) [17, 18] being specially relevant. On the other hand, networks of personal relationships have been inferred from phone calls [19], email exchanges [20] or face-to-face interaction data [21, 22]. When longitudinal data is lacking, network dynamics can be analyzed indirectly using cross-sectional data. The basic idea behind this approach is that the dynamical mechanisms driving the temporal evolution of a network leave a fingerprint in the structure observed at single points in time. This is the key idea behind Exponential Random Graph Models [23] and related methods [24, 25, 26].
This entire corpus of research has led to the conclusion that three main factors drive social network dynamics [27]: (1) attributes of people (people have a number of individual attributes that are not encoded a priori in the network structure in any way and affect the way the network evolves over time, such as shared interests or social traits); (2) endogenous mechanisms (the current structure of the network constrains and modifies the future one, i.e., certain links that appear or disappear by the simple presence of other links in the network), and (3) contextual factors and disruptive events (every network is affected by the environment in which it is embedded, like the institutional setting). These factors operate locally in the network through what we call mechanisms, rules or processes that determine how the structural or functional properties of a network change over time. Some well known mechanisms are: (1) homophily and influence [27], that drive nodes with similar traits to be positively connected; (2) reciprocity [27], that favours the creation of bidirectional links; (3) transitivity and closure [27], that promotes the closing of triangles and the formation of clusters; (4) differential popularity [28], that results in heterogeneous degree distributions shaped by individual traits, levels of activity or sociability, visibility, or preferential attachment; (5) balance [29], that promotes the creation of triangles with an even number of negative links; or (6) resource allocation [2], by which people maintain a finite number of bonds constrained by their resources availability, explaining the volatility of negative links, link removal to free up resources that can be reinvested in other links, or the prevention of the formation of new relationships due to a lack of resources, even if all the conditions for link formation are ideal. This last mechanism will be specially relevant to the appearance of the equilibrium in the network.
@@
Understanding the dynamics of personal relationships is difficult because these mechanisms operate simultaneously, but also because their behaviours are often coupled, giving rise to effects impossible to predict by analysing them separately. Examples of this include the formation of hierarchies and other combinations such as the ones studied in [30], [22],[31],[32], [33] or [28]. It would thus seem that the combination of this myriad of behaviours would make the evolution of the system unpredictable. However, in this paper we show how, despite all this complexity, the temporal evolution of the network can be described by simple mathematical models and, more importantly, its behaviour can be predicted with very high accuracy. Specifically, we provide evidence that relationship networks exhibit a global behaviour akin to equilibrium dynamics. This means that the macroscopic, average properties of the network – such as the distribution of personal contacts, the prevalence of patterns like edges, triangles, or larger structures, and aggregate metrics like network density – remain constant, fluctuating around stable values, even as individual links continuously evolve at a microscopic level. Interestingly, there are apparent trends of stability in other aspects of human social behaviour, which suggest a natural tendency toward the formation of stable patterns and self-reinforcing dynamics. Thus, Hobbs and Burke show how social connections recover after the death of a friend [34], while Alessandreti et al. found that certain mobility patterns are conserved over time [35], linking these patterns to individuals’ social ties. Other studies, such as [36, 37, 38], stress the stability and robustness of online social behaviours. We will circle back to this research in the discussion section.
Despite its importance, to the best of our knowledge, no prior research has explicitly focused on determining whether social networks are indeed in equilibrium. In this context, we present an example of an empirical social network that is in rigorous equilibrium, given by a rich dataset collected over four years. While this is a single case, we will argue in the discussion that the emergence of equilibrium in social networks can be a very general phenomenon. The importance of this result cannot be overstated. If a social network is in equilibrium, its properties can be predicted more accurately over time, facilitating intervention studies. Interventions can be introduced and compared against the equilibrium network, which serves as a baseline for analysis, helping to strengthen social cohesion or address social isolation. The possibility that this equilibrium behaviour may be general also suggests that certain social behaviours and organization principles transcend cultural, situational or individual differences, connecting individual interactions and broader social structures. From a modelling perspective, equilibrium assumptions enable the use of simpler, more tractable mathematical models, such as statistical mechanics tools, developed for physical systems. Conclusions about the mechanisms driving network dynamics, structural trends, and other insights apply not only to a single observed point in time but to the entire unobserved evolution of the network after the transient formation period. This is specially relevant for methods that employ cross-sectional data to investigate network dynamics, like Exponential Random Graph Models. In fact, these approaches rely on an implicit and often overlooked assumption: the network must be in or near equilibrium, as discussed in [39]. In simple terms, when using cross-sectional data, the researcher assumes that every observation is statistically equivalent to any other, a condition that is fulfilled if the network is in equilibrium.
@@ 2 Results
In this work, we analyse a dataset we have collected between 2020 and 2024 that contains the temporal evolution of the network of personal relationships among 888 people belonging to the Blas de Otero High School in Madrid (see the Materials and Methods section for details on the data collection, data composition and data curation). The reported relationships were coded as +2 (very good), +1 (good), -1 (bad), and -2 (very bad). Almost every student provided this list, resulting in the extraction of a weighted directed network for the entire high school, which we will refer to as snapshot or wave. We repeated this process every 16-20 weeks, and with this data we reconstructed the network at different points in time, collecting 10 snapshots of the network over 4 years. With respect to previous literature, we have introduced some improvements in the data collection process. Respondents are allowed to report an unlimited number of relationships, resulting in richer and more heterogeneous data. We also increased the sample size by one order of magnitude, with a larger number of snapshots, allowing for the detailed study of the dynamics in the long term while maintaining a certain level of granularity. Relationships are self-reported, and they are directed and weighted, between -2 and 2, allowing the consideration of negative relationships. Preliminary analyses of earlier, shorter versions of this dataset have been reported elsewhere [40, 41]. It is worth noting that by coding relationships into these four categories (-2, -1, +1 and +2), we inevitably lose many of the details that characterize social interactions. Thus, the resulting network we analyse in this study is not a direct reflection of the complexity of real personal relationships, but rather a simplified representation. Real-world relationships exist on a spectrum, with layers of emotional, social, and contextual subtleties that cannot be completely captured in these discrete categories. However, even with this simplification, we argue that identifying an equilibrium in this network serves as a meaningful proxy for the underlying dynamics of the real network. The equilibrium we detect in this reduced representation suggests that the macroscopic patterns of stability we observe likely reflect deeper robust social structures. If the simplified relationships follow this equilibrium dynamics, it implies that the more complex relationships they represent also tend toward stability over time.
@@
Drawing an analogy to a social system represented as a network, it is essential to define both the macroscopic properties and the microscopic dynamics. In a social network evolving over time, links appear, disappear, or change in nature. From the perspective of a node, an outgoing or incoming +1 link can become a +2, a -1, a -2 or disappear. We refer to an absent link as a 0 link. This constitutes the microscopic dynamics of the system: the evolution of individual ties/links. From these microscopic components, we can build macroscopic properties of the network. In the network, we define an edge as the connection between two nodes that contains two links, one from the first node to the second and one from the second node to the first. Since we have 5 types of links, -2, -1, 0, +1, +2, there are 25 possible edge types. Notice the distinction between link/tie and edge we introduce in our nomenclature. Although it is usual to treat both as interchangeable, we keep this distinction throughout the paper. Our first macroscopic property is the distribution of different edge types: how many +2+2, +1+2, +1+1, etc., edges exist in the network. Notice that a +2+1 edge is not equivalent to a +1+2, specially from a dynamical perspective. Although both edges can evolve towards a +1-1 edge, in the +2+1 case the +2 needs to become a +1 and the +1 a -1, and in the +1+2 case, the +1 remains unchanged and the +2 becomes a -1. Therefore, we keep this distinction in all computations. A second macroscopic property is the in-degree and out-degree distributions of the nodes. For each type of link (+2, +1, -1, and -2), we count how many nodes have 0 incoming +2 links, 1 incoming +2 links, etc. This process is repeated for outgoing links and every type of link, generating eight degree distributions. A third macroscopic property can be the distribution of different triangle types that can form among three nodes. Other macroscopic properties include structural characteristics of the network, such as clustering, average shortest path length, assortativity, etc.
@@
Let us introduce the concept of transition matrix using the edges as our macroscopic property. We define the edges transition matrix $P_{\beta\alpha}$ from snapshot $\alpha$ to snapshot $\beta$ as the matrix whose $ij$ element represents $P(j|i)$, i. e., the conditional probability of an edge of going to the $j$ state in snapshot $\beta$ provided it started in the $i$ state in snapshot $\alpha$ (see Materials and Methods for the construction of this matrix). Additionally, we define the edge state distribution $\pi_{\alpha}$ as a column vector in which each element $i$ is the density of edges of type $i$ in the snapshot $\alpha$. With these two objects, it is direct to see that:
@@
In other words, if we take the distribution of edges in one snapshot and multiply this distribution by the transition matrix, we obtain the distribution of edges in the next snapshot. Since we have 10 snapshots, our data allows us to construct nine transition matrices between consecutive snapshots. The first relevant question is whether these nine transition matrices are statistically equivalent. If they were, it would indicate that the dynamics are stationary, meaning the transition probabilities between edge states remain constant over time. In the Materials and Methods section we explain how to perform such a comparison.
@@
The stationarity of the transition matrices does not necessarily imply that the network is in equilibrium; it may not even have an equilibrium state under these dynamics. As we will see, this is not our case.